OpenStudy (31356):

Polynomial help! (3c^2 d^2)(-5cd^4)^2 MEDAL REWARDED & FAN!!

OpenStudy (31356):

@poopsiedoodle @jim_thompson5910 @phi @rose21

OpenStudy (31356):

@thomaster

OpenStudy (31356):

@Jigglypuff123

OpenStudy (31356):

Any ideas?

OpenStudy (mathmale):

Hello, Friend31356, are you certain that this problem involves a polynomial?

OpenStudy (31356):

Very certain, this HW is titled Pondering Polynomials, so it must be.

OpenStudy (mathmale):

What it says it says, but what it says is not necessarily the truth! I see a negative quantity, but I do not see addition/subtraction of terms.

OpenStudy (31356):

Any ideas how to solve it or steps to get the answer?!

OpenStudy (mathmale):

I do see that "Order of Operations" rules are called for. Familiar with those?

OpenStudy (31356):

Definitely

OpenStudy (mathmale):

To start out solving this problem, we MUST do the exponentiation first. What is (-5cd^4)^2?

OpenStudy (31356):

-10c^2d^8? Is it correct, I am assuming?

OpenStudy (mathmale):

If you like, you could break that down into \[(-5)^{2}c ^{^{2}}(d ^{4})^{2}\]

OpenStudy (mathmale):

Where did that -10 come from?

OpenStudy (31356):

I learned it from Math today, I am positive.

OpenStudy (mathmale):

Remember, Friend 31356, " ^ " represents exponentiation, not multiplication.

OpenStudy (31356):

-5*2=-10

OpenStudy (31356):

Right? @mathmale

OpenStudy (mathmale):

You are multiplying. No Please raise (-5) to the 2nd power (exponentiation).

OpenStudy (31356):

-25?

OpenStudy (mathmale):

\[5^{2}=25;(-5)^{2}=?\]

OpenStudy (31356):

-25

OpenStudy (mathmale):

(-2)(-2)=?

OpenStudy (31356):

4

OpenStudy (mathmale):

(-5)(-5)=?

OpenStudy (31356):

25

OpenStudy (mathmale):

(-5)^2=?

OpenStudy (mathmale):

Bet you know very well that (-5)^2 = (-5)(-5).

OpenStudy (31356):

25

OpenStudy (31356):

Oh. I get it

OpenStudy (31356):

LOL

OpenStudy (mathmale):

Cool. So, Friend, (-5cd^4)^2 = ?

OpenStudy (31356):

25 c^2 d^8

OpenStudy (31356):

Right? @mathmale

OpenStudy (31356):

Thanks for the medal!

OpenStudy (mathmale):

Very nice! Notice that I've awarded you a real medal for good work?

OpenStudy (31356):

Indeed

OpenStudy (31356):

Notice that I've awarded you a real medal for good explanation? LOL

OpenStudy (mathmale):

Indeed. Now please multiply your latest result by 3c^2d^2, and y ou'll then be done. And thank you for the medal!

OpenStudy (31356):

My pleasure

OpenStudy (31356):

I understand the rest now thanks!

OpenStudy (mathmale):

Very happy to work with you. Keep up the good work. Post some more questions! Bye for now.

OpenStudy (31356):

Ok